# From EHRs to Patient Pathways: Scalable Modeling of Longitudinal Health Trajectories with LLMs

### v1

@@ 3.2.3 Length-of-Stay Indicator
Converging to final states like discharge requires accurate state predictions across multiple forecast iterations, which is challenging because such events are relatively rare compared to typical value changes, often causing the model to inadvertently prolong patient stays without termination. We tackle this by appending a Length-of-Stay (LOS) indicator, a simple countdown of remaining hours for each clinical unit $u$. During training we feed a noised ground-truth LOS and ask the model to update it. 50% of the time we drop the indicator entirely so the model learns to infer it from context. When present, we perturb it by ±20% to encourage dynamic adjustments. This approach teaches the model to estimate the LOS, when unavailable during inference and encourages it to re-estimate LOS during simulation rather than rigidly decrementing it.
@@ 3.2.4 Fine-tuning for Outcome Prediction
Table 1: Next Time-Step Prediction: We report event F1, event-value scores for numerical and categorical predictions and the avg./max. context and input tokens required to capture this context.
F1 Numerical $\downarrow$ Categorical Avg./Max. Avg./Max.
macro/micro macro/micro macro/micro context input $\downarrow$
Statistic 0.02 / 0.02 0.99 / 0.97 0.03 / 0.16 n/a n/a
ETHOS 0.003 / 0.03 1.00 / 0.97 0.04 / 0.31 1217 / 2048 1217 / 2048
E2P-T-1h 0.42 / 0.66 0.68 / 0.89 0.32 / 0.55 380 / 3142 380 / 3142
E2P-T-24h 0.46 / 0.80 0.67 / 0.75 0.37 / 0.61 1365 / 4250 1365 / 4250
E2P-S 0.54 / 0.78 0.65 / 0.71 0.41 / 0.62 9823 / 86621 220 / 1494
E2P-T+S 0.44 / 0.79 0.70 / 0.74 0.36 / 0.57 9823 / 86621 220 / 1494
Restricted to data supported by ETHOS:
ETHOS 0.04 / 0.08 0.96 / 0.96 0.86 / 0.85 1217 / 2048 1217 / 2048
E2P-T+S 0.08 / 0.17 0.92 / 0.92 0.99 / 0.88 9823 / 86621 220 / 1494
@@ 4.1 Dataset and Pre-processing
We use the MIMIC-IV database johnson2023mimic , containing de-identified EHRs for approximately 300,000 patients, as it is one of the largest and the most comprehensive EHR database to date, including coded, numerical and free-text features from different clinical units. To demonstrate the generality of our approach, we incorporate nearly all clinically relevant tables, encompassing 22 tables and 20 ICU Chart Event categories (listed in Appendix A), including admissions, lab results, medications, procedures, vital signs, and more, spanning the Emergency Department (ED), hospital wards, and Intensive Care Unit (ICU). To retain real-world complexity, we apply minimal exclusion criteria and preserve missing or incorrectly populated fields. Average context lengths are 380 tokens per hour, 1,880 for 24 hours, or 11,361 for the full history.
Data is split at the patient level: 95% for training, 2.5% each for validation and test sets. Training samples consist of EHR data $D_{p,t}$ up to time $t$, with $t+1$ as the label. In practice, since patient stays often span hundreds of hours, using all hourly slices is computationally infeasible. Instead, we sample time points from all patient based on the following strategies:
Weighted Sampling: Time points are weighted based on the rarity of clinical events in the output, with rarer events receiving higher log-scaled weights.
Oversampling Critical Transitions: Key events such as admissions, discharges, and deaths are over-sampled to ensure adequate representation in training.
Overall, we collect one million weighted samples for training, while fixed sets of 5000 val/test time points each are sampled without weighting, preserving the original event distribution.
@@ 4.2 Longitudinal Simulation Tasks
We define nine clinically relevant evaluation tasks across ED, Hospital, and ICU, focusing on outcome and development prediction, partially inspired by mcdermott2021EHRbenchmark . For each task we perform 500 simulations to assess iterative trajectory prediction. Binary tasks are balanced to ensure robust evaluation of rare adverse events. ED tasks use only ED data, while Hospital and ICU tasks incorporate all prior data. Detailed task descriptions can be found in Appendix E.
Outcome Prediction Tasks: ED Admission (will ED patient be hospitalized), ED/Hospital Discharge Diagnosis (ICD categories), Imminent Mortality (within 24 hours), 3-day Length-of-Stay.
@@ 4.3 Baselines
For comparison, we select trajectory and outcome prediction methods trained on the MIMIC-IV dataset with publicly available code, enabling reproduction on our data splits and tasks:
ETHOS renc2024zero : A transformer-based model for tokenized health timeline simulation, trained on next-token prediction. We compare against ETHOS on next time-step prediction and all applicable trajectory simulation tasks.
MEME lee2024emergency : An LLM-based approach for outcome classification on textualized Emergency Department data, fine-tuned to predict outcomes from the full textual summary of ED records.
REMed pmlr-v252-kim24a : A transformer-based long context model for ICU outcome prediction that scores EHR events using an MLP and feeds the most important ones to a transformer for classification. We compare fine-tuned EHR2Path variants with MEME and REMed on applicable outcome prediction tasks.
@@ 4.4 Evaluation Metrics
Event Prediction: Micro- and macro-averaged F1 scores for predicting which events or features will be present in the next hour.
Event+Value Prediction: Measures overall correctness by requiring both correct event timing and value prediction. Numerical values use a modified MAE (correct-time predictions scored by MAE; incorrect/missing predictions receive maximum error). Numerical values are normalized with 1st–99th percentile clipping and min-max scaling to [0,1]. Categorical values use modified accuracy where missing predictions are counted as incorrect.
@@
Longitudinal Simulation Metrics. For extended trajectory forecasting, we use task-specific metrics. Binary classification tasks (e.g., mortality prediction) are evaluated using accuracy on balanced test sets, while regression-based tasks (e.g., vital sign forecasting) are assessed using event-based F1, MAE on min-max normalized values, and accuracy. Value predictions are only evaluated if a matching time-point exists, with a ±$\pm 1$ hour tolerance to account for EHR recording variability.
@@ 5 Results and Discussion
Event F1 n/a 0.61 0.53 0.57
Value MAE $\downarrow$ n/a 0.10 0.12 0.10
Value Acc. n/a 0.54 0.54 0.51
@@
Event F1 0.05 0.24 0.17 0.23
Value MAE $\downarrow$ 0.21 0.10 0.11 0.08
Value Acc. n/a 0.32 0.36 0.40
@@
Event F1 n/a 0.70 0.75 0.71
Value MAE $\downarrow$ n/a 0.09 0.10 0.08
Value Acc. n/a 0.69 0.80 0.79
@@
bold indicates the best value; underlined the 2nd-best
5.1 Next Timestep Prediction Results
We evaluated our models’ ability to predict the next time-step in a patient’s trajectory, with results in Table 1. We compare our models against a statistical baseline (sampling events by occurrence ratios and using mean/majority values) and ETHOS (renc2024zero, ), a transformer-based timeline simulator. The statistical baseline’s inferior results highlight task complexity. ETHOS under-performs across all metrics, mainly due to its inability to predict many EHR event types. When restricting the evaluation to data supported by ETHOS, ETHOS improves slightly, but E2P-T+S still outperforms, even though it predicts a much wider range of data.
Among our models, E2P-S achieves the highest macro F1 and numerical predictions, demonstrating the effectiveness of our Masked Summarization Bottleneck in capturing essential information from the entire patient history. E2P-T-24h shows strong performance in micro F1, indicating its ability to handle frequent events well and outperforms E2P-T-1h, showing the benefit of including more longitudinal context. The combined model, E2P-T+S, performs comparably to E2P-T-24h but does not surpass E2P-S in overall performance. Additionally, E2P-S and E2P-T+S handle significantly larger contexts, averaging 8$\times$ more, and up to 20$\times$ in MIMIC-IV, while using fewer tokens than text-based models. This shows our summarization mechanism effectively condenses extensive patient histories into manageable embeddings, improving prediction performance and enabling efficient handling of large-scale, heterogeneous EHR data.
5.2 Patient Trajectory Simulation Results
@@
E2P-T-24h, E2P-S, and E2P-T+S all perform strongly on simulation tasks (Table 3), with distinct strengths: E2P-S excels in long-term, dense-context tasks such as ICU Vital Sign and Input Development, while E2P-T-24h outperforms in immediate state predictions such as ED Admission or Diagnosis tasks. The combined model, E2P-T+S consistently ranks first or second, often achieving the best results. This stability makes it a robust compromise, balancing text- and summary-based strengths for strong performance across diverse simulations. A qualitative example of a vital sign simulation of E2P-T+S is shown in Appendix B. Compared to ETHOS, EHR2Path significantly outperforms in the trajectory-based lab value prediction, underscoring our significant improvements in longitudinal EHR modeling, while achieving comparable results in outcome prediction. Further, the comprehensive data integration of EHR2Path enables support for a significantly broader range of prediction tasks. These results highlight our strength in temporal modeling and simulation of patient pathways.
5.3 Fine-tuning for Outcome Prediction
@@
While our primary focus is on trajectory simulation, we also investigate the adaptability of our model as a foundation for outcome prediction tasks through targeted fine-tuning. We report results for Specialized Pathway Fine-Tuning (E2P-FT-P) and Outcome-Oriented Fine-Tuning (E2P-FT-O), and the specialized outcome prediction models MEME (lee2024emergency, ) and REMed (pmlr-v252-kim24a, ) in Table 3. Both fine-tuning strategies demonstrate the model’s ability to effectively specialize for specific outcomes, with the outcome fine-tuning outperforming pathway fine-tuning at the cost of of lacking intermediate outcomes. Compared to the baselines, EHR2Path-FT-O outperforms in three of four tasks. These findings underscore the versatility of our approach beyond trajectory prediction, highlighting its potential as a robust foundation for a wide range of EHR-based prediction tasks.
@@ 5.4 Additional Ablation Results
To assess the impact of key components, we conduct ablation studies on the Length-of-Stay (LOS) Indicator and the Masked Summarization Bottleneck size.
Effect of the LOS Indicator   We assessed the LOS Indicator’s impact on tasks requiring convergence to final states, ED Admission Prediction and Hospital Discharge Diagnosis. Failure to converge within realistic, dataset-derived step limits is counted as an incorrect prediction. As shown in Table 5, removing the LOS Indicator significantly hinders convergence to a final state, reducing predictive accuracy.
Effect of Bottleneck Size   We examine performance across various Masked Summarization Bottleneck sizes, trained for a limited number of steps. While longer contexts add detail, their quadratic cost limits scalability. Table 5 shows that, while larger bottleneck sizes slightly reduce validation loss, the most substantial improvement occurs from 4 to 8 tokens. Based on this, we select a bottleneck size of 8, balancing performance and efficiency.
@@ 5.5 Limitations and Future Work
While the EHR2Path models establish a strong foundation for patient trajectory modeling, the focus on a single healthcare system may limit the generalizability of results to other healthcare settings, with differing demographics and site-specific treatment strategies. Additionally, due to data availability, the current work focuses on short-term pathways and hospital trajectories. While these are highly relevant for clinical decision-making within acute care settings, long-term pathways offer significant promise for capturing broader health trajectories. These limitations could be addressed by collecting and using more comprehensive EHR data from a diverse set of countries and hospitals and by expanding it to include lifetime health data for long-term trajectory modeling. Additionally to increasing the scope of the data, exploring research into robust adaptation techniques across diverse healthcare systems, presents an exciting and impactful research direction for the community.

### v2

@@ 3.2.3 Length-of-Stay Indicator
Converging to final states like discharge requires accurate state predictions across multiple forecast iterations, which is challenging because such events are relatively rare compared to typical value changes, often causing the model to inadvertently prolong patient stays without termination. We tackle this by appending a Length‑of‑Stay (LOS) indicator, a simple countdown of remaining hours for each clinical unit $u$. To increase robustness, and allow the model to work without using the ground-truth LOS token during inference, we perform two augmentations. In half of the samples, we drop the LOS token entirely from the input, and in the other half, we inject noise (+-20%) into the LOS token, encouraging the model to re-estimate the remaining LOS at each time step rather than rigidly decrementing it. Crucially, at inference time, we never include the ground truth LOS token in the input, instead the first step is prompted without LOS token, while in later steps the predicted LOS tokens of prior steps are included.
@@ 4.1 Dataset and Pre-processing
We use the MIMIC-IV database (Johnson et al., 2023b), containing de-identified, real-world EHRs for approximately 300,000 patients, as it is one of the largest and the most comprehensive EHR databases to date, offering granular and heterogeneous patient records, including coded, numerical and free-text features from different clinical units. We incorporate all clinically relevant tables, encompassing 22 tables and 20 ICU Chart Event categories (listed in section˜A.1), including admissions, lab results, medications, procedures, vital signs, and more, spanning the Emergency Department (ED), hospital wards, and Intensive Care Unit (ICU). Further, its inherent noisiness and sparsity strengthen its value as a realistic benchmark. To retain this real-world complexity, we avoid applying exclusion criteria and preserve missing or incorrectly populated fields. Following previous work (Wang et al., 2020; McDermott et al., 2021), we aggregate values hourly using average for numerical values, and most frequent for categorical ones, managing computational feasibility by balancing temporal resolution and recording frequency. Average context lengths are 380 tokens per hour, 1,880 for 24 hours, or 11,361 for the full history. Data is split at the patient level: 95% for training, 2.5% each for validation and test sets. Training samples consist of EHR data $D_{p,t}$ up to time $t$, with $t+1$ as the label. In practice, since patient stays often span hundreds of hours, using all hourly slices is computationally infeasible. Instead, we sample time points from all patients using weighted sampling, where time points are weighted based on the rarity of clinical events in the output, with rarer events receiving higher log-scaled weights. Additionally, we oversample critical transitions and key events such as admissions, discharges, and deaths to ensure adequate representation in training. Overall, we collect one million weighted samples for training, while fixed sets of 5000 val/test time points each are sampled without weighting, preserving the original distribution.
@@ 4.2 Longitudinal Simulation Tasks
We define nine clinically relevant evaluation tasks across ED, Hospital, and ICU, focusing on outcome and development prediction, partially inspired by McDermott et al. (2021). For each task we perform 500 simulations to assess iterative trajectory prediction. ED tasks use only ED data, while Hospital and ICU tasks incorporate all prior data. Detailed task descriptions can be found in section˜A.5.
Outcome Prediction Tasks: ED Admission (will ED patient be hospitalized), ED/Hospital Discharge Diagnosis (multi-label prediction over 18 ICD categories), Imminent Mortality (24 hours), Imminent Discharge (3 days).
@@ 4.3 Baselines
For comparison, we select trajectory and outcome prediction methods trained on the MIMIC-IV dataset with publicly available code, enabling reproduction on our data splits and tasks: ETHOS (Renc et al., 2024), a transformer-based model for tokenized health timeline simulation, trained on next-token prediction. We compare against ETHOS on next time-step prediction and all applicable trajectory simulation tasks. MEME (Lee et al., 2024), an LLM-based approach for outcome classification on textualized Emergency Department data, fine-tuned to predict outcomes from the full textual summary of ED records. REMed (Kim et al., 2024), a transformer-based long context model for ICU outcome prediction that scores EHR events using an MLP and feeds the most important ones to a transformer for classification. We compare fine-tuned EHR2Path variants with MEME and REMed on applicable outcome prediction tasks.
@@ 4.4 Evaluation Metrics
Event Prediction: To accurately represent false positive and negative rates, we use micro- and macro-averaged F1 scores for predicting which events or features will be present in the next hour.
Event+Value Prediction: Measures overall correctness by requiring both correct event timing and value prediction. Numerical values use a modified MAE (correct-time predictions scored by MAE; incorrect/missing predictions receive maximum error). Numerical values are normalized with 1st–99th percentile clipping and min-max scaling to [0,1]. Categorical values use modified accuracy with missing predictions counted as incorrect.
@@
Longitudinal Simulation Metrics. For extended trajectory forecasting, we use task-specific metrics. Binary classification tasks (e.g., mortality prediction) are evaluated using accuracy on balanced test sets, while regression-based tasks (e.g., vital sign forecasting) are assessed using event-based F1, MAE on min-max normalized values, and accuracy. For binary outcome tasks evaluated with accuracy, we use a balanced test set for sample efficiency. Value predictions are only evaluated if a matching time-point exists, with a $\pm 1$ hour tolerance to account for EHR recording variability.
@@ 5 Results and Discussion
We evaluate EHR2Path on MIMIC-IV, jointly modeling emergency department (ED), ward, and ICU pathways as a single longitudinal trajectory. To provide a high-level overview of our evaluation, fig.˜4 summarizes both task coverage and performance across the main simulation and outcome tasks. EHR2Path is evaluated on a substantially broader set of in-hospital pathway tasks than prior baselines, while remaining competitive or superior on directly comparable settings. A detailed analysis of these task groups follows in later sections.
Figure 4: Overview of the evaluation across simulation and outcome tasks, summarizing task coverage and performance of EHR2Path relative to prior baselines on the main task groups. We report the primary metric per task, with numerical tasks reporting $1-\mathrm{MAE}$. Unsupported tasks are plotted as zero to visualize task coverage.
5.1 Next Timestep Prediction Results
Table 1: Next Time-Step Prediction: We report event F1, event-value scores for numerical and categorical predictions and the avg./max. context and input tokens required to capture this context. Unless otherwise specified, E2P-T denotes the 24-hour text-only model. 95% confidence intervals are given in brackets.
F1 Numerical $\downarrow$ Categorical Avg./Max. Avg./Max.
macro/micro macro/micro macro/micro context input
Statistic 0.02 / 0.02 0.99 / 0.97 0.03 / 0.16 n/a n/a
ETHOS 0.003 [-0.25ex](0.00,0.008) / 0.03 [-0.25ex](0.03,0.03) 1.00 [-0.25ex](0.99,1.00) / 0.97 [-0.25ex](0.97,0.97) 0.04 [-0.25ex](0.00,0.12) / 0.31 [-0.25ex](0.31,0.32) 1217/2048 1217/2048
E2P-T-1h 0.42 [-0.25ex](0.30,0.53) / 0.64 [-0.25ex](0.64,0.64) 0.67 [-0.25ex](0.52,0.80) / 0.89 [-0.25ex](0.88,0.89) 0.30 [-0.25ex](0.16,0.44) / 0.52 [-0.25ex](0.51,0.53) 380/3142 380/3142
E2P-T 0.47 [-0.25ex](0.36,0.57) / 0.78 [-0.25ex](0.78,0.78) 0.64 [-0.25ex](0.51,0.77) / 0.74 [-0.25ex](0.74,0.75) 0.33 [-0.25ex](0.21,0.47) / 0.57 [-0.25ex](0.56,0.57) 1365/4250 1365/4250
E2P-S 0.48 [-0.25ex](0.36,0.59) / 0.76 [-0.25ex](0.76,0.77) 0.66 [-0.25ex](0.55,0.78) / 0.72 [-0.25ex](0.71,0.72) 0.33 [-0.25ex](0.21,0.46) / 0.56 [-0.25ex](0.55,0.56) 9823 / 86621 220 / 1494
E2P-S+T 0.43 [-0.25ex](0.33,0.54) / 0.78 [-0.25ex](0.78,0.78) 0.68 [-0.25ex](0.58,0.79) / 0.73 [-0.25ex](0.72,0.73) 0.33 [-0.25ex](0.21,0.46) / 0.55 [-0.25ex](0.55,0.56) 9823/86621 1585/5744
Restricted to data supported by ETHOS:
ETHOS 0.04 [-0.25ex](0.00,0.08) / 0.08 [-0.25ex](0.08,0.09) 0.96 [-0.25ex](0.96,0.96) / 0.96 [-0.25ex](0.96,0.96) 0.86 [-0.25ex](0.86,0.86) / 0.85 [-0.25ex](0.84,0.86) 1217 / 2048 1217 / 2048
E2P-S+T 0.12 [-0.25ex](0.00,0.24) / 0.24 [-0.25ex](0.23,0.25) 0.87 [-0.25ex](0.87,0.87) / 0.87 [-0.25ex](0.87,0.88) 0.99 [-0.25ex](0.99,0.99) / 0.93 [-0.25ex](0.93,0.94) 9823/86621 1585/5744
[…]
5.2 Patient Trajectory Simulation Results
@@
Event F1 n/a 0.61 [-0.25ex](0.58,0.64) 0.53 [-0.25ex](0.49,0.56) 0.57 [-0.25ex](0.54,0.60)
Value MAE $\downarrow$ n/a 0.10 [-0.25ex](0.10,0.10) 0.12 [-0.25ex](0.12,0.13) 0.10 [-0.25ex](0.10,0.11)
Value Acc. n/a 0.54 [-0.25ex](0.48,0.61) 0.54 [-0.25ex](0.46,0.61) 0.51 [-0.25ex](0.44,0.57)
@@
Event F1 0.05 [-0.25ex](0.03,0.07) 0.20 [-0.25ex](0.16,0.25) 0.17 [-0.25ex](0.13,0.20) 0.23 [-0.25ex](0.19,0.27)
Value MAE $\downarrow$ 0.21 [-0.25ex](0.19,0.23) 0.09 [-0.25ex](0.08,0.09) 0.11 [-0.25ex](0.11,0.12) 0.08 [-0.25ex](0.08,0.09)
Value Acc. n/a 0.32 [-0.25ex](0.00,1.00) 0.36 [-0.25ex](0.00,1.00) 0.43 [-0.25ex](0.13,0.81)
@@
Event F1 n/a 0.70 [-0.25ex](0.67,0.72) 0.75 [-0.25ex](0.73,0.78) 0.71 [-0.25ex](0.69,0.73)
Value MAE $\downarrow$ n/a 0.09 [-0.25ex](0.09,0.09) 0.10 [-0.25ex](0.09,0.10) 0.08 [-0.25ex](0.08,0.08)
Value Acc. n/a 0.69 [-0.25ex](0.66,0.72) 0.80 [-0.25ex](0.77,0.82) 0.79 [-0.25ex](0.77,0.81)
@@
E2P-T, E2P-S, and E2P-S+T all perform strongly on simulation tasks (table˜3), with distinct strengths: E2P-S excels in long-term, dense-context tasks such as ICU Vital Sign and Input Development, while E2P-T outperforms in immediate state predictions such as ED Admission or Diagnosis tasks. The combined model, E2P-S+T consistently ranks first or second, often achieving the best results. This stability makes it a robust compromise, balancing text- and summary-based strengths for strong performance across diverse simulations. A qualitative example of a vital sign simulation of E2P-S+T is shown in section˜A.2. Within the subset of shared tasks, our model achieves competitive or superior results compared to ETHOS, and further enables a wide range of tasks beyond its scope due to comprehensive data integration. These results highlight our strength in temporal modeling and simulation of patient pathways.
We further analyze robustness to increasing prediction horizon in fig.˜5. For key development tasks, we gradually increase the horizon in four-hour steps and evaluate event F1 and value MAE at each horizon. As expected, performance decreases with longer horizons, with a more pronounced drop for event occurrence (event F1) than for numerical values. Importantly, the decline is gradual rather than catastrophic, and performance remains at a stable level up to a 24-hour horizon across ED, ward, and ICU.
Figure 5: Development of event detection (F1) and value prediction (MAE) across simulation horizons for ED, hospital, and ICU development tasks. The shaded areas indicate 95% confidence intervals.
5.3 EHR2Path as a Foundation Model for Outcome Prediction
@@
While EHR2Path is designed primarily as a pathway simulator, many clinical applications ultimately require concise outcome predictions: whether a patient will be admitted, what their discharge diagnosis will be, or whether they are at imminent risk of deterioration or discharge from the ICU. We therefore test whether EHR2Path can serve as a general-purpose foundation model for such outcome tasks via task-specific fine-tuning. We report results for Specialized Pathway Fine-Tuning (E2P-FT-P) and Outcome-Oriented Fine-Tuning (E2P-FT-O), and the specialized outcome prediction models MEME (Lee et al., 2024) and REMed (Kim et al., 2024) in table˜3. Both fine-tuning strategies demonstrate the model’s ability to effectively specialize for specific outcomes, with the outcome fine-tuning outperforming pathway fine-tuning at the cost of lacking intermediate outcomes. Compared to the baselines, EHR2Path-FT-O outperforms in three of four tasks. These findings underscore the versatility of our approach beyond trajectory prediction, highlighting its potential as a robust foundation for a wide range of EHR-based prediction tasks.
@@ 5.4 Additional Ablation Results
To assess the impact of key components, we conduct ablation studies on the Length-of-Stay (LOS) Indicator and the Masked Summarization Bottleneck size. We assess the LOS Indicator’s impact on tasks requiring convergence to final states, Hospital Discharge Diagnosis and ICU Imminent Discharge. Failure to converge within realistic, dataset-derived step limits is counted as an incorrect prediction. As shown in table˜5, removing the LOS Indicator significantly hinders convergence to a final state, reducing predictive accuracy. Further, we examine performance across various Masked Summarization Bottleneck sizes, trained for a limited number of steps. While longer contexts add detail, their quadratic cost limits scalability. Table˜5 shows that, while larger bottleneck sizes slightly reduce validation loss, the most substantial improvement occurs from 4 to 8 tokens. Based on this, we select a bottleneck size of 8, balancing performance and efficiency.
@@ 5.5 Limitations and Future Work
While the EHR2Path models establish a strong foundation for patient trajectory modeling, the focus on a single healthcare system may limit the generalizability of results to other healthcare settings, with differing demographics and site-specific treatment strategies, which could be addressed by collecting and using more comprehensive EHR data from a diverse set of countries and hospitals. Further, expanding the approach to include lifetime health data for long-term trajectory modeling is an exciting future direction.